# Benchmarking Post-Training Quantization in LLMs: Comprehensive Taxonomy, Unified Evaluation, and Comparative Analysis

### v1

@@ A.1 Quantization Preliminaries
Quantization aims to reduce inference and storage overheads by converting high precision floating-point values into their corresponding low precision integer counterparts (Nagel et al., 2021). The asymmetric weight-only quantization is formulated as:
@@ A.2 Quantization Surveys and Benchmarks
Most model compression reviews summarize ample concepts, principles, and a coarse-grained classification of commonly used compression methods (Tang et al., 2024; Park et al., 2024; Miao et al., 2023; Wan et al., 2023; Zhu et al., 2023; Yang et al., 2024), but limited focus is placed on quantization. The other literature contributes a more detailed exploration of quantization. Wang et al. (2024) provides comprehensive explanations of quantization frameworks but lacks detailed taxonomy and summaries of their specific characteristics, so that the followers are still confused by how to select a basic framework to further develop. Gong et al. (2024) delves deeper into the existing methods and provides finer distinctions but no experiments are conducted to evaluate the performance of each category. Kurtic et al. (2024) explores the accuracy-performance trade-off for popular quantization formats via broad and automated benchmarks, but only 3 models are included into examination which undermines the persuasiveness of the study. Li et al. (2024) focus on synthesizing experimental results but offers limited valuable guidance for future research. It is evident that, there is still a lack of an technical and development guideline for LLMs PTQ which enables researchers to select correct model for deployment or foundational PTQ strategy based on their requirements. Our benchmark fills this gap by experimentally analyzing the characteristics of different PTQ strategies and providing recommendations for researchers to consider.
Figure 3: The number of quantization papers since the explosion of LLMs (2022).
@@ B.1 Detailed Experimental Setup
AWQ (Lin et al., 2024), GPTQ (Frantar et al., 2022), OmniQuant (Shao et al., 2023), and QuIP (Chee et al., 2024) serve as representatives of the four PTQ strategies in our taxonomy, owing to their superior performance and broad practical deployment in numerous released LLMs. All baselines perform channel-wise quantization and set up as their description.
@@
To demonstrate the universality and generalizability of our benchmark, we select LLaMA family, the most influential and widely used open-source LLM family, for performance evaluation. Specifically, this includes LLaMA-1 (7B to 70B) (Touvron et al., 2023a), LLaMA-2 (7B to 70B) (Touvron et al., 2023b), as well as LLaMA-3 and LLaMA-3.1 (8B and 70B) (Dubey et al., 2024), covering a substantial range of model sizes. Unlike other benchmarks, we further broaden our scope to investigate the cross-modality and cross-architecture capabilities of different PTQ strategies. Particularly, Mixtral (Jiang et al., 2024), DeepSeeK-MoE (Dai et al., 2024) and Mamba (Gu and Dao, 2023) are used to evaluate cross-architecture capabilities. LLaVA-1.5 (Liu et al., 2024a) and VILA-1.5 (Lin et al., 2023) are employed to assess cross-modality capabilities.
@@
For LLMs, MoE LLMs and Mamba, we evaluate the perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020), and zero-shot accuracies on 9 commonsense reasoning tasks, including WinoGrande (Sakaguchi et al., 2021), Race (Lai et al., 2017), LAMBADA (Paperno et al., 2016), PIQA (Bisk et al., 2020), MMLU (Hendrycks et al., 2020), CEval (Huang et al., 2023), GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019) and ARC (Clark et al., 2018). For MLLMs, we evaluate their visual-language reasoning abilities on GQA (Hudson and Manning, 2019), VizWiz (Gurari et al., 2018), ScienceQA (Lu et al., 2022), TextVQA (Singh et al., 2019), POPE (Li et al., 2023b), MMBench (Liu et al., 2025) and MMVet (Yu et al., 2023).
@@
For all evaluated models, the calibration data consists of 128 random 2048 token-segments from WikiText2. Using a batch size of 1, the entire quantization and inference processes are implemented using PyTorch (Paszke et al., 2019) package and deployed on 8 NVIDIA A800-80G GPUs. Codes for evaluation will be packed and open-sourced in the future.
@@ B.2 Detailed Evaluation Results
In this section, we present the detailed perplexity and reasoning accuracy on specific tasks. Table 7 and Table 8 show the detailed results on undertrained LLMs (correspond to Table 1). Table 9 and Table 10 shows the detailed results on fully trained LLMs (correspond to Table 2). Table 11 shows the detailed results on MoE LLMs (correspond to Table 4). Table 12 shows the detailed results on Mamba (correspond to Table 3). Table 13 shows the detailed results on Multimodal LLMs (correspond to Table 5).
@@
Table 7: Detailed performance of the four baselines on LLaMA models.
@@
Mixtral-$8\times 7$B-w4 GPTQ 75.85 40.19 76.36 71.20 81.66 65.96 44.73 54.06 62.08 83.16 55.20 4.18 7.11
OmniQ 75.45 40.38 76.36 70.35 81.72 65.46 44.80 53.53 63.15 83.33 54.95 4.18 7.20
Mixtral-$8\times 7$B-w3 GPTQ 72.38 39.52 72.46 64.23 80.03 59.66 34.55 35.63 60.53 80.26 50.17 4.99 7.86
OmniQ 71.51 39.81 67.28 58.78 81.34 57.20 38.04 24.56 58.50 79.17 49.57 5.20 8.62
Mixtral-$8\times 7$B-w2 GPTQ 51.85 28.42 13.14 8.67 60.34 24.35 26.00 0.15 34.02 42.55 20.56 21.02 24.81
OmniQ 49.72 22.49 0.00 0.00 52.18 22.97 23.03 0.00 25.96 26.68 21.33 3.4e3 6.0e3
DeepSeekMoE-16B-w4 GPTQ 68.59 38.85 72.06 67.81 79.54 37.12 31.43 14.56 57.56 76.26 44.03 6.65 9.25
OmniQ 68.82 39.14 68.60 62.80 78.35 36.45 34.70 12.96 56.56 74.71 42.83 6.85 9.57
DeepSeekMoE-16B-w3 GPTQ 67.88 38.28 66.89 60.80 78.56 33.42 29.87 8.87 55.10 73.86 40.27 7.35 10.14
OmniQ 65.43 35.69 59.56 51.99 77.04 24.87 25.85 5.23 53.99 72.05 37.29 8.01 11.23
DeepSeekMoE-16B-w2 GPTQ 53.20 28.61 10.63 1.26 62.73 24.75 26.52 0.30 33.30 40.28 23.12 55.45 42.75
OmniQ 50.67 25.45 2.33 2.33 60.83 23.11 23.11 0.15 30.53 36.15 20.73 72.41 78.40
@@
Mamba-1.4B-w4 GPTQ 59.35 33.30 64.06 55.79 73.39 25.24 22.96 2.05 44.19 64.35 29.61 11.24 14.24
OmniQ 51.93 28.13 2.99 1.42 59.19 22.92 26.15 0.00 29.08 36.28 24.15 1.1e4 7.3e4
@@
Mamba-1.4B-w3 GPTQ 57.06 32.25 53.66 45.95 71.93 26.79 22.96 0.99 41.61 60.06 28.67 13.96 17.31
OmniQ 49.88 26.79 7.86 9.80 59.19 24.90 26.15 0.00 31.85 40.19 24.32 1.7e3 1.6e3
@@
Mamba-1.4B-w2 GPTQ 51.38 24.50 2.64 1.11 58.05 23.93 22.66 0.00 29.04 35.23 21.33 1.2e3 577.89
OmniQ 48.78 24.69 0.00 0.00 54.24 23.22 25.78 0.00 26.33 31.36 19.11 3.6e4 2.3e4
@@
Mamba-2.8B-w4 GPTQ 64.17 32.63 67.53 60.64 74.65 25.56 24.74 1.29 48.53 68.86 33.96 9.88 12.71
OmniQ 60.30 32.92 43.26 40.42 69.31 23.58 22.88 0.99 41.28 58.96 27.22 18.38 21.94
@@
Mamba-2.8B-w3 GPTQ 62.59 33.49 63.24 55.00 73.72 25.76 23.18 0.76 46.75 65.91 31.83 11.92 14.78
OmniQ 56.91 32.15 26.59 21.68 68.01 26.19 25.04 0.15 40.91 56.36 30.29 35.21 38.62
@@
Mamba-2.8B-w2 GPTQ 49.25 25.26 0.37 0.37 57.34 25.22 23.25 0.00 28.45 35.23 19.88 854.82 537.60
OmniQ 48.78 22.78 0.64 0.29 56.04 24.85 23.40 0.00 28.47 29.97 20.31 1.7e4 6.7e3
@@
Table 12: Detailed performance of the four baselines on Mamba.
Models Methods GQA Vizwiz ScienceQA TextVQA POPE MMBench MMBench-cn MM-Vet Avg.
LLaVA1.5-7B-w4 GPTQ 61.80 50.79 69.25 57.58 86.90 64.26 55.41 31.40 61.74
OmniQ 61.44 49.28 68.12 56.30 87.4 62.89 54.73 33.50 61.27
QuIP 61.65 49.77 68.30 57.43 87.20 62.47 54.07 33.00 61.32
AWQ 61.73 49.27 69.75 57.67 87.00 64.18 56.70 33.20 61.97
LLaVA1.5-7B-w3 GPTQ 60.64 52.11 67.72 55.11 84.30 62.03 47.25 28.30 57.18
OmniQ 58.07 52.54 62.70 51.98 81.30 59.71 42.78 27.20 54.54
QuIP 57.51 53.07 62.61 49.71 87.50 46.31 44.60 29.80 53.89
AWQ 60.48 53.61 66.73 55.97 85.80 64.95 52.66 29.20 58.68
LLaVA1.5-7B-w2 GPTQ 24.65 0.08 1.96 13.65 68.60 0.95 0.69 4.60 14.40
OmniQ 0.00 0.00 0.75 0.87 68.00 1.29 1.63 4.80 9.67
QuIP 30.83 19.76 20.77 18.45 71.40 4.81 1.80 8.10 21.99
AWQ 0.00 0.00 0.00 0.00 68.00 0.00 0.00 0.50 8.56
LLaVA1.5-13B-w3 GPTQ 62.70 48.67 72.46 59.68 87.00 67.87 60.91 35.00 63.75
OmniQ 62.01 50.82 70.55 58.23 89.20 64.52 54.73 32.60 62.38
QuIP 62.60 54.45 71.87 59.70 87.90 67.87 58.85 33.70 64.02
AWQ 62.69 52.99 71.78 59.84 88.50 67.78 60.05 34.30 64.08
LLaVA1.5-13B-w2 GPTQ 47.52 37.27 26.88 35.98 88.20 23.54 2.66 16.90 38.46
OmniQ 0.00 0.00 0.00 0.00 68.00 1.46 0.77 0.00 7.80
QuIP 56.25 50.30 38.65 41.93 76.00 44.76 11.25 18.80 45.59
AWQ 0.00 0.00 0.00 0.00 68.00 0.60 0.43 0.90 7.77
VILA1.5-8B-w4 GPTQ 33.20 61.91 78.52 62.78 55.30 70.62 57.22 39.30 54.64
OmniQ 33.16 60.03 69.11 60.83 45.80 56.01 50.52 42.00 49.92
QuIP 30.58 61.01 73.80 58.28 72.40 59.54 55.76 38.60 53.43
AWQ 33.02 60.40 73.94 62.03 45.40 65.29 56.70 41.30 51.99
VILA1.5-8B-w3 GPTQ 33.20 55.31 68.69 55.37 55.30 70.62 48.88 34.80 52.77
OmniQ 33.16 51.70 52.13 52.82 45.80 56.01 10.48 31.50 41.70
QuIP 30.58 58.07 67.27 59.31 72.40 59.54 48.02 39.10 54.29
AWQ 33.02 58.70 73.38 62.32 45.40 65.29 46.48 37.60 52.77
VILA1.5-8B-w2 GPTQ 0.15 0.98 9.22 1.11 68.00 66.70 0.17 3.30 18.70
OmniQ 0.00 0.01 0.05 0.27 68.00 66.70 0.95 4.30 17.54
QuIP 4.83 16.17 11.72 1.07 62.00 60.20 1.89 14.10 21.50
AWQ 0.00 0.00 0.00 0.00 68.00 66.70 0.00 0.50 16.90
VILA1.5-13B-w3 GPTQ 63.98 62.32 81.58 65.17 86.80 74.31 63.66 43.90 69.34
OmniQ 60.35 62.37 78.07 65.00 89.90 70.88 61.94 40.00 67.68
QuIP 62.98 63.44 80.92 66.12 86.90 73.02 63.23 41.80 68.96
AWQ 63.17 61.75 81.75 66.29 89.20 73.97 66.24 44.50 69.91
VILA1.5-13B-w2 GPTQ 51.34 45.70 49.78 44.54 80.30 38.66 2.66 27.70 45.93
OmniQ 0.00 0.10 0.21 1.75 67.90 1.03 0.26 4.10 8.39
QuIP 55.70 46.53 63.17 50.16 82.60 56.01 25.43 31.70 54.18
AWQ 0.00 0.00 0.00 0.00 68.00 0.52 0.69 0.00 7.69

### v4

@@ A.1 Quantization Preliminaries
Quantization aims to reduce inference and storage overheads by converting high precision floating-point values into their corresponding low precision integer counterparts [52]. The asymmetric weight-only quantization is formulated as:
@@ A.2 Quantization Surveys and Benchmarks
Most model compression reviews summarize ample concepts, principles, and a coarse-grained classification of commonly used compression methods [64, 55, 51, 68, 80, 72], but limited focus is placed on quantization. The other literature contributes a more detailed exploration of quantization. Wang et al. [70] provides comprehensive explanations of quantization frameworks but lacks detailed taxonomy and summaries of their specific characteristics, so that the followers are still confused by how to select a basic framework to further develop. Gong et al. [22] delves deeper into the existing methods and provides finer distinctions but no experiments are conducted to evaluate the performance of each category. Kurtic et al. [34] explores the accuracy-performance trade-off for popular quantization formats via broad and automated benchmarks, but only 3 models are included into examination which undermines the persuasiveness of the study. Li et al. [39] focus on synthesizing experimental results but offers limited valuable guidance for future research. It is evident that, there is still a lack of an technical and development guideline for LLMs PTQ which enables researchers to select correct model for deployment or foundational PTQ strategy based on their requirements. Our benchmark fills this gap by experimentally analyzing the characteristics of different PTQ strategies and providing recommendations for researchers to consider.
Figure 3: The number of quantization papers since 2022.
@@ B.1 Detailed Experimental Setup
AWQ [42], GPTQ [20], OmniQuant [61], and QuIP [7] serve as representatives of the four PTQ strategies in our taxonomy, owing to their superior performance and broad practical deployment in numerous released LLMs. All baselines perform channel-wise quantization and set up as their description.
@@
To demonstrate the universality and generalizability of our benchmark, we select LLaMA family, the most influential and widely used open-source LLM family, for performance evaluation. Specifically, this includes LLaMA-1 (7B to 70B) [65], LLaMA-2 (7B to 70B) [66], as well as LLaMA-3 and LLaMA-3.1 (8B and 70B) [16], covering a substantial range of model sizes. Unlike other benchmarks, we further broaden our scope to investigate the cross-modality and cross-structure capabilities of different PTQ strategies. Particularly, Mixtral [32], DeepSeeK-MoE [12] and Mamba [24] are used to evaluate cross-structure capabilities. LLaVA-1.5 [43] and VILA-1.5 [41] are employed to assess cross-modality capabilities.
@@
For LLMs, MoE LLMs and Mamba, we evaluate the perplexity on WikiText2 [50] and C4 [57], and zero-shot accuracies on 9 commonsense reasoning tasks, including WinoGrande [58], Race [35], LAMBADA [54], PIQA [6], MMLU [28], CEval [30], GSM8K [11], HellaSwag [77] and ARC [10]. For MLLMs, we evaluate their visual-language reasoning abilities on VQA [3], GQA [31], VizWiz [26], ScienceQA [48], TextVQA [63], POPE [40], MMBench [45] and MMVet [74].
@@
For all evaluated models, the calibration data consists of 128 random 2048 token-segments from WikiText2. Using a batch size of 1, the entire quantization and inference processes are implemented using PyTorch [56] package and deployed on 8 NVIDIA A800-80G GPUs. Codes for evaluation will be packed and open-sourced in the future.
@@ B.2 Detailed Evaluation Results
In this section, we present the detailed perplexity and reasoning accuracy on specific tasks. Table 6, 7, 8 and 9 show the detailed results on LLaMA families (correspond to Table 1 and 2). Table 10 and 11 show the detailed results on MoE and Mamba LLMs (correspond to Table 3). Table 12 shows the detailed results on Multimodal LLMs (correspond to Table 4). Although our experimental evaluation is comprehensive and covers all major strategies of existing weight-only PTQ frameworks, the scale of the experiments limits us to evaluating only the most representative algorithms within each category and several latest baselines. We regard this as a limitation of this paper and will be fixed in our future work.
B.3 Full-precision Results of LLMs
In this section, we present the detailed full-precision perplexity and reasoning accuracy on specific tasks of all the evaluated LLMs. Table 13 shows the results on LLaMA families, MoE LLMs and Mamba LLMs. Table 14 demonstrates the results on VILA1.5 and LLaVA1.5 families.
Appendix C Analysis on Certain Scenarios
C.1 Why Salience-based Strategy Collapses at 2-bit?
AWQ employs scaling hyper-parameters to adjust salient weight channels, but the scaling hyper-parameters are too small. Given that 2-bit quantization restricts weights to only 4 discrete values, most scaled salient weights will be ultimately mapped to the same quantization grid as their unscaled counterparts, making 2-bit AWQ almost degenerate into 2-bit naive RTN (round to nearest).
C.2 Why Optimization-based Strategy Collapses for LLaMA-3/3.1?
Compared with other strategies, OmniQuant’s optimization mechanism is particularly vulnerable to quantization error accumulation due to its reliance on quantized outputs from preceding layers as supervisory information for current-layer optimization. In fully trained LLMs like LLaMA-3/3.1, their pretrained weights carry higher information density so quantization will bring more negative impacts on them [33, 2]. This triggers a cascading effect: severe error accumulation degrades current-layer output validity, which in turn corrupts the quantization supervision for subsequent layers, ultimately creating a detrimental feedback loop.
C.3 Why Compensation-based and Rotaion-based Strategies Demonstrate Robustness?
We summarize that compensation-based method (GPTQ) and rotation-based method (QuIP) demosntrate superior robustness at cross-structure, low-bit and cross training-level scenario in Section 3. Our analysis about this is as follows:
Cross-structure robustness
Unlike AWQ and OmniQuant, the pipeline of both GPTQ and QuIP don’t depends on previous layers (e.g. fuse some hyper-parameters into previous layers) or quantized outputs from preceding layers, instead they make full concentration on the weights in current layer and full-precision outputs from preceding layers.
Low-bit and cross training-level robustness
GPTQ achieves real-time dynamic compensation during its column-wise quantization process, where the more quantization error current weight column generates, the more compensation will be added to the remaining unquantized weights. This error-adaptive mechanism significantly reduces the overall quantization error across the entire weight matrix. QuIP employs random orthogonal matrices to preprocess model weights to enhance incoherence among weights so that the mutual interference post-quantization will be reduced. In addition, this preprocessing induces more uniform magnitude distributions along quantization directions, and the improved uniformity minimizes rounding errors during quantization, ultimately mitigating the performance degradation caused by quantization.
Table 6: Detailed performance of the four baselines on LLaMA models.
@@ C.3 Why Compensation-based and Rotaion-based Strategies Demonstrate Robustness?
Mixtral-$8\times 7$B GPTQ 75.85 40.19 76.36 71.20 81.66 65.96 44.73 54.06 62.08 83.16 55.20 4.18 7.11
-w4 OmniQ 75.45 40.38 76.36 70.35 81.72 65.46 44.80 53.53 63.15 83.33 54.95 4.18 7.20
Mixtral-$8\times 7$B GPTQ 72.38 39.52 72.46 64.23 80.03 59.66 34.55 35.63 60.53 80.26 50.17 4.99 7.86
-w3 OmniQ 71.51 39.81 67.28 58.78 81.34 57.20 38.04 24.56 58.50 79.17 49.57 5.20 8.62
Mixtral-$8\times 7$B GPTQ 51.85 28.42 13.14 8.67 60.34 24.35 26.00 0.15 34.02 42.55 20.56 21.02 24.81
-w2 OmniQ 49.72 22.49 0.00 0.00 52.18 22.97 23.03 0.00 25.96 26.68 21.33 3.4e3 6.0e3
DeepSeekMoE GPTQ 68.59 38.85 72.06 67.81 79.54 37.12 31.43 14.56 57.56 76.26 44.03 6.65 9.25
-16B-w4 OmniQ 68.82 39.14 68.60 62.80 78.35 36.45 34.70 12.96 56.56 74.71 42.83 6.85 9.57
DeepSeekMoE GPTQ 67.88 38.28 66.89 60.80 78.56 33.42 29.87 8.87 55.10 73.86 40.27 7.35 10.14
-16B-w3 OmniQ 65.43 35.69 59.56 51.99 77.04 24.87 25.85 5.23 53.99 72.05 37.29 8.01 11.23
DeepSeekMoE GPTQ 53.20 28.61 10.63 1.26 62.73 24.75 26.52 0.30 33.30 40.28 23.12 55.45 42.75
-16B-w2 OmniQ 50.67 25.45 2.33 2.33 60.83 23.11 23.11 0.15 30.53 36.15 20.73 72.41 78.40
Table 11: Detailed performance of the four baselines on Mamba.
Mamba-1.4B GPTQ 59.35 33.30 64.06 55.79 73.39 25.24 22.96 2.05 44.19 64.35 29.61 11.24 14.24
-w4 OmniQ 51.93 28.13 2.99 1.42 59.19 22.92 26.15 0.00 29.08 36.28 24.15 1.1e4 7.3e4
@@
Mamba-1.4B GPTQ 57.06 32.25 53.66 45.95 71.93 26.79 22.96 0.99 41.61 60.06 28.67 13.96 17.31
-w3 OmniQ 49.88 26.79 7.86 9.80 59.19 24.90 26.15 0.00 31.85 40.19 24.32 1.7e3 1.6e3
@@
Mamba-1.4B GPTQ 51.38 24.50 2.64 1.11 58.05 23.93 22.66 0.00 29.04 35.23 21.33 1.2e3 577.89
-w2 OmniQ 48.78 24.69 0.00 0.00 54.24 23.22 25.78 0.00 26.33 31.36 19.11 3.6e4 2.3e4
@@
Mamba-2.8B GPTQ 64.17 32.63 67.53 60.64 74.65 25.56 24.74 1.29 48.53 68.86 33.96 9.88 12.71
-w4 OmniQ 60.30 32.92 43.26 40.42 69.31 23.58 22.88 0.99 41.28 58.96 27.22 18.38 21.94
@@
Mamba-2.8B GPTQ 62.59 33.49 63.24 55.00 73.72 25.76 23.18 0.76 46.75 65.91 31.83 11.92 14.78
-w3 OmniQ 56.91 32.15 26.59 21.68 68.01 26.19 25.04 0.15 40.91 56.36 30.29 35.21 38.62
@@
Mamba-2.8B GPTQ 49.25 25.26 0.37 0.37 57.34 25.22 23.25 0.00 28.45 35.23 19.88 854.82 537.60
-w2 OmniQ 48.78 22.78 0.64 0.29 56.04 24.85 23.40 0.00 28.47 29.97 20.31 1.7e4 6.7e3
@@
Models Methods VQAv2 GQA Vizwiz ScienceQA TextVQA POPE MMBench MMBench-cn MM-Vet Avg.
LLaVA1.5-7B-w4 GPTQ 78.30 61.80 50.79 69.25 57.58 86.90 64.26 55.41 31.40 61.74
OmniQ 77.80 61.44 49.28 68.12 56.30 87.40 62.89 54.73 33.50 61.27
QuIP 78.00 61.65 49.77 68.30 57.43 87.20 62.47 54.07 33.00 61.32
AWQ 78.20 61.73 49.27 69.75 57.67 87.00 64.18 56.70 33.20 61.97
LLaVA1.5-7B-w3 GPTQ 77.06 60.64 52.11 67.72 55.11 84.30 62.03 47.25 28.30 59.36
OmniQ 76.21 58.07 52.54 62.70 51.98 81.30 59.71 42.78 27.20 56.94
QuIP 74.50 57.51 53.07 62.61 49.71 87.50 46.31 44.60 29.80 56.18
AWQ 77.43 60.48 53.61 66.73 55.97 85.80 64.95 52.66 29.20 60.76
LLaVA1.5-7B-w2 GPTQ 41.89 24.65 0.08 1.96 13.65 68.60 0.95 0.69 4.60 17.45
OmniQ 0.11 0.00 0.00 0.75 0.87 68.00 1.29 1.63 4.80 8.61
QuIP 47.26 30.83 19.76 20.77 18.45 71.40 4.81 1.80 8.10 24.80
AWQ 0.00 0.00 0.00 0.00 0.00 68.00 0.00 0.00 0.50 7.61
LLaVA1.5-13B-w3 GPTQ 79.50 62.70 48.67 72.46 59.68 87.00 67.87 60.91 35.00 63.75
OmniQ 78.80 62.01 50.82 70.55 58.23 89.20 64.52 54.73 32.60 62.38
QuIP 79.20 62.60 54.45 71.87 59.70 87.90 67.87 58.85 33.70 64.02
AWQ 78.80 62.69 52.99 71.78 59.84 88.50 67.78 60.05 34.30 64.08
LLaVA1.5-13B-w2 GPTQ 67.15 47.52 37.27 26.88 35.98 88.20 23.54 2.66 16.90 38.46
OmniQ 0.01 0.00 0.00 0.00 0.00 68.00 1.46 0.77 0.00 7.80
QuIP 72.40 56.25 50.30 38.65 41.93 76.00 44.76 11.25 18.80 45.59
AWQ 0.00 0.00 0.00 0.00 0.00 68.00 0.60 0.43 0.90 7.77
VILA1.5-7B-w4 GPTQ 82.18 63.00 64.22 78.76 66.01 87.50 71.39 63.32 40.90 68.59
OmniQ 81.91 62.95 62.80 78.14 65.69 86.80 70.96 61.77 38.40 67.71
QuIP 81.22 61.53 56.56 73.38 63.09 87.40 66.15 45.96 36.40 63.52
AWQ 82.17 63.26 63.87 79.01 65.95 87.80 71.65 63.23 40.00 68.55
VILA1.5-7B-w3 GPTQ 81.10 60.93 61.53 76.35 64.48 83.90 68.56 60.22 36.50 65.95
OmniQ 80.88 61.58 59.83 74.53 63.27 86.10 67.01 56.62 38.50 65.37
QuIP 77.11 53.50 47.34 51.29 53.35 81.40 45.19 5.93 36.00 50.12
AWQ 81.38 62.09 62.06 76.75 64.06 85.80 68.04 58.33 38.20 66.30
VILA1.5-7B-w2 GPTQ 32.81 18.66 4.32 28.20 13.32 66.10 2.23 1.46 6.60 19.30
OmniQ 40.97 15.88 9.51 2.00 16.09 74.90 4.98 0.34 15.20 19.99
QuIP 60.13 38.20 23.80 25.47 20.48 79.20 0.86 0.17 16.50 29.42
AWQ 0.00 0.00 0.00 0.00 0.00 68.00 0.00 0.00 0.00 7.56
VILA1.5-13B-w3 GPTQ 82.30 63.98 62.32 81.58 65.17 86.80 74.31 63.66 43.90 69.33
OmniQ 80.60 60.35 62.37 78.07 65.00 89.90 70.88 61.94 40.00 67.68
QuIP 82.20 62.98 63.44 80.92 66.12 86.90 73.02 63.23 41.80 68.96
AWQ 82.30 63.17 61.75 81.75 66.29 89.20 73.97 66.24 44.50 69.91
VILA1.5-13B-w2 GPTQ 72.70 51.34 45.70 49.78 44.54 80.30 38.66 2.66 27.70 45.93
OmniQ 0.15 0.00 0.10 0.21 1.75 67.90 1.03 0.26 4.10 8.39
QuIP 76.30 55.70 46.53 63.17 50.16 82.60 56.01 25.43 31.70 54.18
AWQ 0.00 0.00 0.00 0.00 0.00 68.00 0.52 0.69 0.00 7.69
Table 13: Detailed FP16 performance of all evaluated LLMs, MoE LLMs and Mamba LLMs.
Models WinoG Race Lamb-o Lamb-s PiQA MMLU CEval GSM8K HellaS ARC-e ARC-c Wiki C4
LLaMA-7B 70.01 40.29 73.57 67.82 78.67 32.18 26.89 9.02 56.99 75.29 41.81 5.68 7.08
LLaMA-13B 72.77 39.62 76.15 71.08 79.16 43.48 24.44 17.13 59.92 77.36 46.42 5.09 6.61
LLaMA-30B 75.69 40.57 77.59 73.34 80.96 54.63 32.10 33.06 63.34 80.39 52.82 4.10 5.98
LLaMA-65B 77.35 41.53 79.10 74.89 81.34 59.37 38.93 46.85 64.57 81.36 52.82 3.53 5.62
LLaMA2-7B 69.06 39.52 73.86 68.23 78.07 41.84 29.87 13.57 57.14 76.30 43.34 5.47 6.97
LLaMA2-13B 72.14 40.57 76.77 70.33 79.11 52.10 36.85 23.12 60.04 79.46 48.46 4.88 6.47
LLaMA2-70B 77.90 42.58 79.58 74.69 82.21 65.46 44.28 53.30 64.79 82.70 54.35 3.31 5.52
LLaMA3-8B 73.24 40.29 75.65 68.72 79.54 62.23 48.44 50.87 60.14 80.09 50.17 6.14 8.88
LLaMA3-70B 80.66 42.01 79.33 73.94 82.32 75.20 64.56 80.14 66.29 86.78 60.41 2.86 6.73
LLaMA3.1-8B 73.95 39.04 75.45 67.13 80.14 63.35 49.03 49.28 59.99 81.40 51.54 6.27 8.99
LLaMAa3.1-70B 79.72 40.19 78.75 73.86 82.86 75.32 63.37 80.52 66.44 87.21 60.92 2.83 6.70
Mixtral-8x7B 76.48 39.90 78.01 73.10 82.43 68.00 48.66 58.76 64.79 84.01 56.74 3.84 6.88
DeepSeekMoE-16B 70.09 39.23 72.97 69.36 78.89 37.79 35.81 16.45 58.04 76.05 44.54 6.51 9.05
Mamba-130m 52.49 27.37 44.23 33.86 64.53 22.67 23.03 0.76 30.81 47.98 19.80 20.66 22.82
Mamba-370m 55.17 29.67 55.58 47.25 69.48 22.89 22.59 0.61 37.19 55.05 24.91 14.34 17.28
Mamba-790m 56.12 33.01 61.42 54.18 72.09 23.77 23.85 1.29 42.31 61.15 26.45 12.04 14.94
Mamba-1.4B 61.40 34.35 64.91 56.72 74.10 24.87 22.88 0.99 45.03 65.49 29.78 10.76 13.61
Mamba-2.8B 63.14 33.78 69.09 61.32 75.19 26.45 23.77 1.74 49.49 69.78 34.39 9.46 12.28
Table 14: Detailed FP16 performance of all evaluated MLLMs.
Models VQAv2 GQA Vizwiz ScienceQA TextVQA POPE-r POPE-r POPE-a MMBench MMBench-cn MM-Vet
LLaVA1.5-7B 78.52 61.94 50.06 70.22 58.21 87.30 86.10 84.20 64.69 58.08 30.90
LLaVA1.5-13B 80.00 63.25 53.61 74.94 61.19 87.10 86.30 84.50 68.47 63.49 36.60
VILA1.5-7B 82.31 63.39 63.91 79.42 66.10 87.50 86.30 84.90 72.16 63.92 40.30
VILA1.5-13B 82.80 64.38 62.67 83.38 64.97 87.50 86.20 85.20 75.00 66.32 44.20